# Distribution-Aligned Decoding for Efficient LLM Task Adaptation

### v1

@@ Abstract
Adapting billion-parameter language models to a downstream task is still costly, even with parameter-efficient fine-tuning (PEFT). We re-cast task adaptation as output-distribution alignment: the objective is to steer the output distribution toward the task distribution directly during decoding rather than indirectly through weight updates. Building on this view, we introduce Steering Vector Decoding (SVD), a lightweight, PEFT-compatible, and theoretically grounded method. We start with a short warm-start fine-tune and extract a task-aware steering vector from the Kullback-Leibler (KL) divergence gradient between the output distribution of the warm-started and pre-trained models. This steering vector is then used to guide the decoding process to steer the model’s output distribution towards the task distribution. We theoretically prove that SVD is first-order equivalent to the gradient step of full fine-tuning and derive a globally optimal solution for the strength of the steering vector. Across three tasks and nine benchmarks, SVD paired with four standard PEFT methods improves multiple-choice accuracy by up to 5 points and open-ended truthfulness by 2 points, with similar gains (1-2 points) on commonsense datasets without adding trainable parameters beyond the PEFT adapter. SVD thus offers a lightweight, theoretically grounded path to stronger task adaptation for large language models.
@@ 1 Introdcution
Steering Vector Decoding (SVD). To achieve this goal, we present Steering Vector Decoding (SVD), an innovative, efficient, and PEFT-compatible method for task adaptation. SVD begins with a short warm-start fine-tuning phase to obtain the warm-started model, whose output distribution is closer to the task-specific target compared to the pre-trained model. Then we can capture the task-specific direction from the differences between output distributions of the warm-started model and the pre-trained model. Specifically, we first compute the KL divergence between these two distributions, and then use the negative gradient of the KL divergence to construct the steering signal. Next, this signal is mapped from the distribution space to the logit space to avoid simplex geometry violation and yields a task-aware steering vector that tells us which tokens need more (or less) probability mass and by how much. Additionally, confidence-aware constraints are applied to the steering vector to ensure its robustness and stability. Finally, the steering vector is used to adjust the model’s logits at each step during decoding, effectively steering the generation process towards the desired task behavior. Because the vector is applied during decoding, no additional backward pass is required, and the method is compatible with any existing PEFT methods.
Our contributions are summarized as follows: 1) We rethink LLM task adaptation from the perspective of output distribution alignment. 2) We propose the SVD method, which leverages negative gradients of KL divergence between distributions to construct task-aware steering vectors for decoding-time adaptation. 3) We provide theoretical analysis, linking SVD to traditional PEFT methods and derive an analytical solution for the optimal steering strength. 4) We demonstrate through extensive experiments on various tasks and models that SVD, when combined with standard PEFT techniques, consistently improves performance while maintaining computational efficiency.
@@ 3.1 Steering Vector Construction
Figure 1: Illustration of the framework of our proposed SVD. It includes two steps: (a) steering vector construction and (b) task-aware steering vector decoding. After the decoding with the steering vector, we can see the warm-started model’s output distribution is steered towards the task-specific target distribution, thereby enhancing the performance of the model on the downstream task.
@@ 4.1 Experimental Setup
Table 2: Experimental results on commonsense reasoning tasks. We evaluate different PEFT methods and our proposed SVD method on Qwen2.5-7B and LLaMA3.1-8B.
@@ 4.2 Main Results
Multiple-Choice Tasks. Table 1 shows that our approach consistently outperforms baseline PEFT methods. For Qwen2.5-1.5B, SVD with LoRA improves scores from 46.99% to 49.77%. For Qwen2.5-7B, SVD with Prompt Tuning increases scores from 45.49% to 50.29%. For LLaMA3.1-8B, SVD with LoRA boosts scores from 42.89% to 47.80%. Despite occasional MC1 score drops, MC2 and MC3 improvements ensure overall better performance, highlighting SVD’s effectiveness in enhancing truthful answer selection.
Open-Ended Generation Tasks. Table 1 shows that our approach improves performance across all datasets. For Qwen2.5-1.5B, SVD with LoRA increases the score from 44.71% to 45.73%. For Qwen2.5-7B, SVD with P-Tuning v2 raises the score from 62.80% to 64.43%. For LLaMA3.1-8B, SVD with LoRA boosts the score from 45.78% to 46.54%. This demonstrates SVD’s effectiveness in enhancing model responses.
@@
Table 4: Integrating SVD with four basic decoding strategies, Greedy Search, Beam Search, Top-p sampling, and Top-k sampling, where Beam-4 indicates using 4 beams. We evaluate our proposed SVD method on Qwen2.5-7B.
@@ 4.3 Ablation Study
Integrated With Different Basic Decoding Strategies. Table 4 presents experimental results for commonsense reasoning tasks, examining the integration of SVD with various decoding strategies including Greedy Search, Beam Search, Top-p sampling, and Top-k sampling. The results demonstrate that incorporating SVD consistently enhances the performance of fine-tuned LLMs on commonsense reasoning datasets, irrespective of the underlying decoding approach used.
@@ 5 Conclusion
In this paper, we have re-framed LLM task adaptation as a problem of output-distribution alignment. Building on this perspective, we have introduced Steering Vector Decoding (SVD), a lightweight, PEFT-compatible method that can consistently improve performance across a wide range of tasks and model sizes via adjusting the decoding distribution by the task-specific steering vector with a global optimal steering strength. In addition, we have proved the equivalence between SVD and the gradient step of fine-tuning, thereby grounding the method in the classical optimization theory. In summary, SVD offers a lightweight, theoretically grounded, and empirically validated path toward stronger LLM task adaptation, bridging the gap between gradient-based fine-tuning and decoding-time control of model behavior, and demonstrating that shifting distributions, not weights, can be the shortest route to better performance.
@@ 6 Acknowledgement
The research work described in this paper was conducted in the JC STEM Lab of Smart City funded by The Hong Kong Jockey Club Charities Trust under Contract 2023-0108. The work was supported in part by the Hong Kong SAR Government under the Global STEM Professorship and Research Talent Hub. The work of S. Hu was supported in part by the Hong Kong Innovation and Technology Commission under InnoHK Project CIMDA.
@@ A.2 LLM Task Adaptation
Researchers have developed diverse approaches to adapt pre-trained LLMs for downstream tasks [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] by investigating optimal configurations of pre-training data, model architecture, and weight selection. For instance, Cui et al. [48] leverage Earth Mover’s Distance to select the top $K$ most relevant classes from the source domain for task-specific pre-training. Yoon et al. [49] apply reinforcement learning techniques to determine appropriate weights for source domain classes. Other approaches assess model transferability to target domains by examining inter-class covariance relationships between source data and target classes, or by analyzing conditional cross-entropy between source and target labels [50, 51]. More recent research focuses on identifying which pre-trained model weights to fine-tune while keeping others frozen [11, 14, 52]. Zhang et al. [11] introduced gradient-based parameter selection (GPS), a method that uses gradients to identify optimal tunable weights. Fu et al. [52] developed a second-order approximation method (SAM) that approximates the Hessian matrix in the loss function’s second-order Taylor expansion to enhance weight selection precision.
@@ A.3 LLM Decoding Strategies
Appendix B Mathematical Analysis: Equivalence Between SVD and Fine-Tuning
In this appendix, we prove that a first-order Steering-Vector-Decoding (SVD) step is equivalent, in expectation, to one parameter-update step of maximum-likelihood fine-tuning.
@@ B.2 Fine-Tuning Gradient in Logit Space
B.3 One SVD Step in Logit Space
SVD perturbs logits during decoding by
@@ B.4 First-Order Equivalence Theorem
Let $x$ be fixed and let $p_{\phi}(\cdot\mid x)$ be the warm-started distribution. For any label distribution $q(\cdot)$ and any strength $\mu$, denote the KL divergence after one SVD step by
@@
In particular, if $q=\hat{p}_{\text{task}}$ and $p_{\phi}$ is obtained by any fine-tuning algorithm that has converged to a stationary point $(\nabla_{z_{\phi}}\!\mathrm{KL}(q\|p_{\phi})=\mathbf{0})$, then $\partial\mathcal{K}(\mu)/\partial\mu|_{\mu=0}=0$. Hence an infinitesimal SVD step leaves the fine-tuning objective unchanged up to $O(\mu^{2})$.
@@
Eq. 31 states that the first-order of the KL objective reacts to an SVD step exactly as it reacts to a gradient step of fine-tuning. Therefore SVD and fine-tuning are locally equivalent in the space of output distributions, even though one edits logits at decode time while the other edits weights during training.
@@ B.5 Conditions for Higher-Order Equivalence
Therefore, SVD can be viewed as an on-the-fly proxy for one gradient step of fine-tuning, executed in logit space with provable first-order equivalence but without the memory or time overhead of back-propagation. This theoretical link explains why combining SVD with any PEFT method consistently improves performance while preserving efficiency.
@@ D.2 From Tokens to a Global Constant $\bar{\mu}$
This scalar $\bar{\mu}$ not only captures the dominant shift dictated by the task distribution but also preserves the key advantage of SVD: no run-time optimisation loop and no per-token label needed.
@@ Appendix E Experiment Implementation Details
E.1 Implementation Details of SVD
In this section, we provide the implementation details of the SVD method. It is summarized in the following Table 6.
Table 6: Implementation Details of SVD.
@@ E.6 Details about DeepSeek-V3-0324 Evaluation
Table 11: More experimental results on commonsense reasoning tasks. We evaluate different PEFT methods and our proposed SVD method on LLaMA2-7B
@@ F.1 More Results on Commonsense Reasoning Tasks
We conducted additional experiments on commonsense reasoning tasks using the LLaMA2-7B model, comparing various PEFT methods with and without the integration of our proposed SVD method. As shown in Table 11, the results indicate that the SVD-enhanced versions consistently outperform their counterparts across all tasks. Specifically, the average accuracy improvements with SVD are notable: LoRA improves from 31.24% to 33.12%, IA3 from 49.64% to 50.54%, Prompt Tuning from 37.67% to 38.99%, and P-Tuning v2 from 32.10% to 33.50%. These findings underscore the effectiveness of the SVD method in enhancing model performance on commonsense reasoning tasks.
@@ F.2 Comparison with Other Decoding Adaptation Methods
To investigate whether SVD is more beneficial to adapte LLMs on downstream tasks, we expanded our evaluation by comparing SVD with other decoding adaptation techniques, such as TaD [62]. The experimental results, as shown in Table 12, clearly demonstrate that the integration of SVD substantially improves model performance. These findings highlight the critical contribution of SVD to effectively optimizing model capabilities for downstream applications.
Table 12: Comparing SVD with other decoding adaptation techniques. We evaluate our proposed SVD method and TaD method on Qwen2.5-7B.
@@ Appendix G Limitations and Future Work
A primary limitation of Steering Vector Decoding (SVD) is its dependency on an initial warm-start fine-tuning phase to identify an effective, task-specific steering direction. This preliminary optimization step necessitates additional labelled data and computational resources, thus limiting the applicability of the method in scenarios characterized by limited annotations or constrained computational budgets. Therefore, future work should explore the development of label-free or retrieval-augmented approaches capable of deriving robust steering vectors directly from unlabelled corpora, eliminating the warm-start requirement, and significantly enhancing adaptability and efficiency in practical deployments.
@@ Appendix H Practical Impact
Steering Vector Decoding (SVD) transforms task adaptation from a heavyweight weight-update problem into a lightweight distribution-alignment procedure executed entirely at decode time. Below we outline the concrete benefits that make SVD immediately useful in production and research deployments of LLMs.
@@
    Deployment-time efficiency. SVD requires warm start to extract a task-specific steering direction and thereafter operates without further backward passes, optimizer states, or gradient checkpoints. Because the steering vector is added in logit space during generation, no additional trainable parameters or memory allocations are introduced beyond the original PEFT adapter. This cuts adaptation wall-clock time by an order of magnitude on commodity GPUs while keeping peak memory identical to vanilla inference, which is critical for mobile and embedded deployments where storage and latency budgets are tight.
@@
    Consistent accuracy gains at negligible cost. Across three tasks and nine benchmarks, pairing SVD with four standard PEFT methods lifts multiple-choice accuracy by up to 5 percentage points and open-ended truthfulness by 2 percentage points, and adds a 1–2 percentage points average boost on eight commonsense-reasoning datasets. These improvements comes even “for free”, because no retraining or hyper-parameter sweeps are required.
@@
    Plug-and-play compatibility. Because SVD perturbs logits rather than weights, it can be stacked on any PEFT recipe (LoRA, IA3, Prompt Tuning, P-Tuning v2) and on any decoding strategy.
@@
    Theoretically grounded. The SVD step is provably equivalent to the gradient step of maximum-likelihood fine-tuning. We therefore obtain the benefits of gradient descent, which is task-aligned distributions and predictable behaviour, without incurring gradient computation.
By turning task adaptation into a constant-overhead inference-time operation, SVD lowers the barrier to customised LLM deployment for small labs, edge devices, and fast-changing domains where rapid iteration is crucial. Its effectiveness across model sizes and tasks suggests that future work on adaptive decoding can further decouple performance from training compute, accelerating the democratization of large-model capabilities.

### v3

@@ Abstract
Adapting billion-parameter language models to a downstream task is still costly, even with parameter-efficient fine-tuning (PEFT). We re-cast task adaptation as output-distribution alignment: the objective is to steer the output distribution toward the task distribution directly during decoding rather than indirectly through weight updates. Building on this view, we introduce Steering Vector Decoding (SVDecode), a lightweight, PEFT-compatible, and theoretically grounded method. We start with a short warm-start fine-tune and extract a task-aware steering vector from the Kullback-Leibler (KL) divergence gradient between the output distribution of the warm-started and pre-trained models. This steering vector is then used to guide the decoding process to steer the model’s output distribution towards the task distribution. We theoretically prove that SVDecode is first-order equivalent to the gradient step of full fine-tuning and derive a globally optimal solution for the strength of the steering vector. Across three tasks and nine benchmarks, SVDecode paired with four standard PEFT methods improves multiple-choice accuracy by up to 5 percentage points and open-ended truthfulness by 2 percentage points, with similar gains (1-2 percentage points) on commonsense datasets without adding trainable parameters beyond the PEFT adapter. SVDecode thus offers a lightweight, theoretically grounded path to stronger task adaptation for large language models.
@@ 1 Introdcution
Steering Vector Decoding (SVDecode). To achieve this goal, we present Steering Vector Decoding (SVDecode), an innovative, efficient, and PEFT-compatible method for task adaptation. SVDecode begins with a short warm-start fine-tuning phase to obtain the warm-started model, whose output distribution is closer to the task-specific target compared to the pre-trained model. Then we can capture the task-specific direction from the differences between output distributions of the warm-started model and the pre-trained model. Specifically, we first compute the KL divergence between these two distributions, and then use the negative gradient of the KL divergence to construct the steering signal. Next, this signal is mapped from the distribution space to the logit space to avoid simplex geometry violation and yields a task-aware steering vector that tells us which tokens need more (or less) probability mass and by how much. Additionally, confidence-aware constraints are applied to the steering vector to ensure its robustness and stability. Finally, the steering vector is used to adjust the model’s logits at each step during decoding, effectively steering the generation process towards the desired task behavior. Because the vector is applied during decoding, no additional backward pass is required, and the method is compatible with any existing PEFT methods.
Our contributions are summarized as follows: 1) We rethink LLM task adaptation from the perspective of output distribution alignment. 2) We propose the SVDecode method, which leverages negative gradients of KL divergence between distributions to construct task-aware steering vectors for decoding-time adaptation. 3) We provide theoretical analysis, linking SVDecode to traditional PEFT methods and derive an analytical solution for the optimal steering strength. 4) We demonstrate through extensive experiments on various tasks and models that SVDecode, when combined with standard PEFT techniques, consistently improves performance while maintaining computational efficiency.
@@ 3.1 Steering Vector Construction
Figure 1: Illustration of the framework of our proposed SVDecode. It includes two steps: (a) steering vector construction and (b) task-aware steering vector decoding. After the decoding with the steering vector, we can see the warm-started model’s output distribution is steered towards the task-specific target distribution, thereby enhancing the performance of the model on the downstream task.
@@ 4.1 Experimental Setup
Table 2: Experimental results on commonsense reasoning tasks. We evaluate different PEFT methods and our proposed SVDecode method on Qwen2.5-7B and LLaMA3.1-8B.
@@ 4.2 Main Results
Multiple-Choice Tasks. Table 1 shows that our approach consistently outperforms baseline PEFT methods. For Qwen2.5-1.5B, SVDecode with LoRA improves scores from 46.99% to 49.77%. For Qwen2.5-7B, SVDecode with Prompt Tuning increases scores from 45.49% to 50.29%. For LLaMA3.1-8B, SVDecode with LoRA boosts scores from 42.89% to 47.80%. Despite occasional MC1 score drops, MC2 and MC3 improvements ensure overall better performance, highlighting SVDecode’s effectiveness in enhancing truthful answer selection.
Qwen2.5-7B %Truth %Info %T*I
w 55.48 50.61 46.95
w/o 0.02 0.01 0.00
Open-Ended Generation Tasks. Table 1 shows that our approach improves performance across all datasets. For Qwen2.5-1.5B, SVDecode with LoRA increases the score from 44.71% to 45.73%. For Qwen2.5-7B, SVDecode with P-Tuning v2 raises the score from 62.80% to 64.43%. For LLaMA3.1-8B, SVDecode with LoRA boosts the score from 45.78% to 46.54%. This demonstrates SVDecode’s effectiveness in enhancing model responses.
@@
Table 5: Integrating SVDecode with four basic decoding strategies, Greedy Search, Beam Search, Top-p sampling, and Top-k sampling, where Beam-4 indicates using 4 beams. We evaluate our proposed SVDecode method on Qwen2.5-7B.
@@ 4.3 Ablation Study
Integrated With Different Basic Decoding Strategies. Table 5 presents experimental results for commonsense reasoning tasks, examining the integration of SVDecode with various decoding strategies including Greedy Search, Beam Search, Top-p sampling, and Top-k sampling. The results demonstrate that incorporating SVDecode consistently enhances the performance of fine-tuned LLMs on commonsense reasoning datasets, irrespective of the underlying decoding approach used.
@@ 5 Conclusion
In this paper, we have re-framed LLM task adaptation as a problem of output-distribution alignment. Building on this perspective, we have introduced Steering Vector Decoding (SVDecode), a lightweight, PEFT-compatible method that can consistently improve performance across a wide range of tasks and model sizes via adjusting the decoding distribution by the task-specific steering vector with a global optimal steering strength. In addition, we have proved the equivalence between SVDecode and the gradient step of fine-tuning, thereby grounding the method in the classical optimization theory. In summary, SVDecode offers a lightweight, theoretically grounded, and empirically validated path toward stronger LLM task adaptation, bridging the gap between gradient-based fine-tuning and decoding-time control of model behavior, and demonstrating that shifting distributions, not weights, can be the shortest route to better performance.
@@ 6 Acknowledgement
The research work described in this paper was conducted in the JC STEM Lab of Smart City funded by The Hong Kong Jockey Club Charities Trust under Contract 2023-0108. The work was supported in part by the Hong Kong SAR Government under the Global STEM Professorship and Research Talent Hub. The work of Senkang Hu was supported in part by the Hong Kong Innovation and Technology Commission under InnoHK Project CIMDA.
@@ A.2 LLM Task Adaptation
Researchers have developed diverse approaches to adapt pre-trained LLMs for downstream tasks by investigating optimal configurations of pre-training data, model architecture, and weight selection. For instance, Cui et al. [37] leverage Earth Mover’s Distance to select the top $K$ most relevant classes from the source domain for task-specific pre-training. Yoon et al. [38] apply reinforcement learning techniques to determine appropriate weights for source domain classes. Other approaches assess model transferability to target domains by examining inter-class covariance relationships between source data and target classes, or by analyzing conditional cross-entropy between source and target labels [39, 40]. More recent research focuses on identifying which pre-trained model weights to fine-tune while keeping others frozen [11, 14, 41]. Zhang et al. [11] introduced gradient-based parameter selection (GPS), a method that uses gradients to identify optimal tunable weights. Fu et al. [41] developed a second-order approximation method (SAM) that approximates the Hessian matrix in the loss function’s second-order Taylor expansion to enhance weight selection precision.
@@ A.3 LLM Decoding Strategies
Appendix B Mathematical Analysis: Equivalence Between SVDecode and Fine-Tuning
In this appendix, we prove that a first-order Steering-Vector-Decoding (SVDecode) step is equivalent, in expectation, to one parameter-update step of maximum-likelihood fine-tuning.
@@ B.2 Fine-Tuning Gradient in Logit Space
B.3 One SVDecode Step in Logit Space
SVDecode perturbs logits during decoding by
@@ B.4 First-Order Equivalence Theorem
Let $x$ be fixed and let $p_{\phi}(\cdot\mid x)$ be the warm-started distribution. For any label distribution $q(\cdot)$ and any strength $\mu$, denote the KL divergence after one SVDecode step by
@@
In particular, if $q=\hat{p}_{\text{task}}$ and $p_{\phi}$ is obtained by any fine-tuning algorithm that has converged to a stationary point $(\nabla_{z_{\phi}}\!\mathrm{KL}(q\|p_{\phi})=\mathbf{0})$, then $\partial\mathcal{K}(\mu)/\partial\mu|_{\mu=0}=0$. Hence an infinitesimal SVDecode step leaves the fine-tuning objective unchanged up to $O(\mu^{2})$.
@@
Eq. 31 states that the first-order of the KL objective reacts to an SVDecode step exactly as it reacts to a gradient step of fine-tuning. Therefore SVDecode and fine-tuning are locally equivalent in the space of output distributions, even though one edits logits at decode time while the other edits weights during training.
@@ B.5 Conditions for Higher-Order Equivalence
Therefore, SVDecode can be viewed as an on-the-fly proxy for one gradient step of fine-tuning, executed in logit space with provable first-order equivalence but without the memory or time overhead of back-propagation. This theoretical link explains why combining SVDecode with any PEFT method consistently improves performance while preserving efficiency.
@@ D.2 From Tokens to a Global Constant $\bar{\mu}$
This scalar $\bar{\mu}$ not only captures the dominant shift dictated by the task distribution but also preserves the key advantage of SVDecode: no run-time optimisation loop and no per-token label needed.
@@ Appendix E Experiment Implementation Details
E.1 Implementation Details of SVDecode
In this section, we provide the implementation details of the SVDecode method. It is summarized in the following Table 6.
Table 6: Implementation Details of SVDecode.
@@ E.6 Details about DeepSeek-V3-0324 Evaluation
Table 11: More experimental results on commonsense reasoning tasks. We evaluate different PEFT methods and our proposed SVDecode method on LLaMA2-7B
@@ F.1 More Results on Commonsense Reasoning Tasks
We conducted additional experiments on commonsense reasoning tasks using the LLaMA2-7B model, comparing various PEFT methods with and without the integration of our proposed SVDecode method. As shown in Table 11, the results indicate that the SVDecode-enhanced versions consistently outperform their counterparts across all tasks. Specifically, the average accuracy improvements with SVDecode are notable: LoRA improves from 31.24% to 33.12%, IA3 from 49.64% to 50.54%, Prompt Tuning from 37.67% to 38.99%, and P-Tuning v2 from 32.10% to 33.50%. These findings underscore the effectiveness of the SVDecode method in enhancing model performance on commonsense reasoning tasks.
@@ F.2 Comparison with Other Decoding Adaptation Methods
To investigate whether SVDecode is more beneficial to adapte LLMs on downstream tasks, we expanded our evaluation by comparing SVDecode with other decoding adaptation techniques, such as TaD [51]. The experimental results, as shown in Table 12, clearly demonstrate that the integration of SVDecode substantially improves model performance. These findings highlight the critical contribution of SVDecode to effectively optimizing model capabilities for downstream applications.
Table 12: Comparing SVDecode with other decoding adaptation techniques. We evaluate our proposed SVDecode method and TaD method on Qwen2.5-7B.
@@ Appendix G Limitations and Future Work
A primary limitation of Steering Vector Decoding (SVDecode) is its dependency on an initial warm-start fine-tuning phase to identify an effective, task-specific steering direction. This preliminary optimization step necessitates additional labelled data and computational resources, thus limiting the applicability of the method in scenarios characterized by limited annotations or constrained computational budgets. Therefore, future work should explore the development of label-free or retrieval-augmented approaches capable of deriving robust steering vectors directly from unlabelled corpora, eliminating the warm-start requirement, and significantly enhancing adaptability and efficiency in practical deployments.
@@ Appendix H Practical Impact
Steering Vector Decoding (SVDecode) transforms task adaptation from a heavyweight weight-update problem into a lightweight distribution-alignment procedure executed entirely at decode time. Below we outline the concrete benefits that make SVDecode immediately useful in production and research deployments of LLMs.
@@
    Deployment-time efficiency. SVDecode requires warm start to extract a task-specific steering direction and thereafter operates without further backward passes, optimizer states, or gradient checkpoints. Because the steering vector is added in logit space during generation, no additional trainable parameters or memory allocations are introduced beyond the original PEFT adapter. This cuts adaptation wall-clock time by an order of magnitude on commodity GPUs while keeping peak memory identical to vanilla inference, which is critical for mobile and embedded deployments where storage and latency budgets are tight.
@@
    Consistent accuracy gains at negligible cost. Across three tasks and nine benchmarks, pairing SVDecode with four standard PEFT methods lifts multiple-choice accuracy by up to 5 percentage points and open-ended truthfulness by 2 percentage points, and adds a 1–2 percentage points average boost on eight commonsense-reasoning datasets. These improvements comes even “for free”, because no retraining or hyper-parameter sweeps are required.
@@
    Plug-and-play compatibility. Because SVDecode perturbs logits rather than weights, it can be stacked on any PEFT recipe (LoRA, IA3, Prompt Tuning, P-Tuning v2) and on any decoding strategy.
@@
    Theoretically grounded. The SVDecode step is provably equivalent to the gradient step of maximum-likelihood fine-tuning. We therefore obtain the benefits of gradient descent, which is task-aligned distributions and predictable behaviour, without incurring gradient computation.
By turning task adaptation into a constant-overhead inference-time operation, SVDecode lowers the barrier to customised LLM deployment for small labs, edge devices, and fast-changing domains where rapid iteration is crucial. Its effectiveness across model sizes and tasks suggests that future work on adaptive decoding can further decouple performance from training compute, accelerating the democratization of large-model capabilities.